# GenCode: A Generic Data Augmentation Framework for Boosting Deep Learning-Based Code Understanding

### v2

@@ Abstract
Pre-trained code models lead the era of code intelligence with multiple models have been designed with impressive performance. However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in this field. In this paper, we introduce a generic data augmentation framework, GenCode, to enhance the training of code understanding models. Simply speaking, GenCode follows a generation-and-selection paradigm to prepare useful training code data. Specifically, it employs code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics. To evaluate the effectiveness of GenCode, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5). Compared to the state-of-the-art (SOTA) code augmentation method, MixCode, GenCode produces code models with 2.92% higher accuracy and 4.90% robustness on average.
Introduction
@@
Deep learning (DL) for code has been one of the successful DL applications over the years. The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection (Zhou et al. 2019), is commonly known as code learning. From the model level, different DL models specifically designed for code learning have been proposed, such as CodeBERT (Feng et al. 2020), GraphCodeBERT (Guo et al. 2020), and CodeT5 (Wang et al. 2021). From the commercial perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development. As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  (Lu et al. 2021) and program classification (Puri et al. 2021), DL already achieved significantly better results than traditional SE techniques.
@@
However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models (Sun et al. 2022). The challenge behind this is preparing training data for code models is heavy work. Although today developers can easily collect raw code data from the open-source community such as GitHub  [^3] https://github.com, data labeling is still expensive, especially for code-related tasks. It needs huge human effort with professional domain knowledge.
@@
To tackle this problem, data augmentation is a famous technique to increase the diversity and quality of training data based only on existing labeled data. In the code learning field techniques such as code refactoring methods and text-oriented transformation methods can be directly used for code augmentation. Unfortunately, existing studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022) stated that code refactoring methods have limited advantages in the accuracy improvement of code models. Besides,  (Dong et al. 2023b) revealed that code models trained using existing data augmentation methods still suffer from the low robustness issue. Moreover, a recent work, MixCode (Dong et al. 2023a), that follows the spirit of Mixup (Zhang et al. 2018) to do code augmentation achieved promising results, but it can be only used for classification tasks. Therefore, there are currently no effective and reliable code augmentation methods that can be generalized across different tasks.
@@
To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision (Cubuk et al. 2020; Yang et al. 2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding. Specifically, at each training epoch, GenCode first generates training candidates by using different code transformation methods. Here, we consider different types of methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods. After that, GenCode selects valuable training data from the candidates with the guidance of importance metrics. The important metric used in GenCode is simply defined as the loss value of code data produced by the code model. Finally, the selected samples are used for the model training. We conducted comprehensive experiments to study how GenCode can improve the accuracy and robustness of code models. The experimental results demonstrated that, compared to the SOTA method, MixCode, GenCode can produce code models with a 2.92% higher accuracy and 4.90% better robustness on average.
@@
  • We propose the first generation-and-selection-based framework, GenCode, which automatically enriches the training data for boosting the code understanding models.
@@
  • We empirically demonstrate that GenCode can produce more accurate and robust code understanding models compared to existing data augmentation methods and conduct an ablation study to explore how different importance scores affect the effectiveness of GenCode.
Background and Related Work
Code Representation
@@
Code representation is a crucial technique that converts source code into a format interpretable by Deep Neural Networks (DNNs), facilitating effective learning of source code features. Since source code can be seen as sequences of tokens or structural graph (Allamanis et al. 2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated program classification (Puri et al. 2021) and code clone detection (Wang et al. 2020). Existing methods for code representation can be roughly divided into two groups, sequence-based representation and graph-based representation. Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens (Alon et al. 2019). Graph-based representation learns the structural information of code. Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information (Dinella et al. 2020; Yasunaga and Liang 2020).
@@
Unlike code representation, our work focuses on code data preprocessing and studies the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
Data Augmentation for Code Understanding
@@
Data augmentation is a well-studied technique in the machine learning field that increases the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort. For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym (Feng et al. 2021).
@@
Recently, data augmentation has gained attention from software engineering researchers. Various data augmentation methods designed for source code have been proposed (Gao et al. 2020; Yefet, Alon, and Yahav 2020; Allamanis, Jackson-Flux, and Brockschmidt 2021; Pour et al. 2021; Bui, Yu, and Jiang 2021; Wang et al. 2022; Park, Kim, and Han 2023; Yang et al. 2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming and dead store, to create more code data without breaking the semantics of the original codes. Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information (Lacerda et al. 2020).
@@
Existing works (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; de Paula Rodrigues, Braga, and Dahab 2023; Dong et al. 2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models. For instance, a recent empirical study (Dong et al. 2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation. Besides, the SOTA code data augmentation method, MixCode, which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring. However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
@@
Compared to existing works, we propose a simple yet effective code augmentation framework that is not only unaffected by the dataset and model constraints but also applicable to both classification and non-classification tasks.
Definition
@@
Definition 1.
A program code  [^4] For simplicity, we use code in the paper. is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
Definition 2.
A code model is defined as a 3-tuple $M=\left(parameter,\mathcal{P},Y\right)$, where $parameter$ contains all the weights in code model $M$, $\mathcal{P}$ is the input code, and $Y$ is the label of $\mathcal{P}$ given a specific downstream code task.
Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 (Wang et al. 2021).
Definition 3.
Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)\&\mathcal{P^{\prime}}.content\neq\mathcal{P}.content$.
Here, code transformation methods can be any method that can change the content of program statements such as text data transformation methods (e.g., synonym replacement and code refactoring methods (e.g., dead if statement adding).
Definition 4.
Given a code model $M$, a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, semantic-preserving (breaking) code augmentation indicates $\mathcal{P}.se=\mathcal{P^{\prime}}.se$ $M(\mathcal{P})=M(\mathcal{P^{\prime}})$ ($\mathcal{P}.se\neq\mathcal{P^{\prime}}.se$ $M(\mathcal{P})\neq M(\mathcal{P^{\prime}})$).
Definition 5.
@@
Assumption and Preliminary Study
GenCode selects the most valuable data for the code model training. This approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training. To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data. Specifically, the study is conducted on the Google Code Jam (GCJ) (Yang et al. 2022) dataset and CodeBERT pre-trained code model. The detailed steps are as follows:
@@
Step 4. We use the Pearson Correlation Coefficient (PCC) (Schober, Boer, and Schwarte 2018) metric to check the correlation between the loss values and the accuracies.
Table 1: Correlation between loss values and code model accuracy. The P-value less than 0.01 is highlighted with a blue background.
Coefficient P-value
Initialized (epoch 0) 0.6075 7.9687e-12
Middle (epoch 10) 0.2440 0.0125
Well-trained (epoch 20) 0.1657 0.0926
Table 1 summarizes the results of our preliminary study. Here, we considered three types of code models, the initialed model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios. The results indicate that there is a positive correlation between the loss values of data and the accuracy of code models trained by using these data. From the initialized model to the well-trained model, this correlation becomes weaker but still exists. We conjecture that when the model is in the early stage, most of the data that have high loss values are useful for the model training. However, when training proceeds, the data that still have high loss values are the model itself cannot learn. Thus, there will be a correlation drop from the early epoch to the later ones. Based on this phenomenon, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
@@
Methodology
Overview
The goal of GenCode is to produce code models with better performance on test data by training models on carefully generated data. Given a training set of program $\mathcal{P}_{train}$, its transformed version $\mathcal{P^{\prime}}_{train}$, and a set of test programs $\mathcal{P}_{test}$, after the training process $\mathsf{train}\left(\cdot\right)$ , $\varrho(M^{\prime}(\mathcal{P}_{test}))>\varrho(M(\mathcal{P}_{test}))$, where $M.parameter=\mathsf{train}(M,\mathcal{P}_{train})$ $M^{\prime}.parameter=\mathsf{train}(M^{\prime},\mathcal{P^{\prime}}_{train})$, and $\varrho\left(\cdot\right)$ is the performance measurement.
GenCode follows the generation-and-selection workflow and utilizes different types of code transformation methods to generate diverse code data first, then selects code data that are most useful for training as the training data. Algorithm 1 shows how GenCode works. In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14). Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data. Overall, GenCode contains two main steps, utilizing code data augmentation methods to compose the search space and importance-guided data selection.

### v3

@@ Abstract
Pre-trained code models lead the era of code intelligence, with multiple models designed with impressive performance. However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in this field. In this paper, we introduce a generic data augmentation framework, GenCode, to enhance the training of code understanding models. Simply speaking, GenCode follows a generation-and-selection paradigm to prepare useful training code data. Specifically, it employs code augmentation techniques to generate new code candidates first and then identifies important ones as the training data by influence scores. To evaluate the effectiveness of GenCode, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5) and two recent released code-specific Large Language Models (LLMs) (e.g., Qwen2.5-Coder). Compared to the state-of-the-art (SOTA) code augmentation method MixCode, GenCode produces pre-trained code models with 2.92% higher accuracy and 4.90% adversarial robustness on average. For code-specific LLMs, GenCode achieves an average improvement of 0.93% in accuracy and 0.98% in natural robustness.
1 Introduction
@@
Deep learning (DL) for code has been one of the successful DL applications over the years. The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection Zhou et al. (2019), is commonly known as code learning. From the model level, different DL models specifically designed for code learning have been proposed, such as CodeBERT Feng et al. (2020), CodeT5 Wang et al. (2021), and Qwen2.5-Coder Hui et al. (2024). From the commercial perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development. As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  Lu et al. (2021) and program classification Puri et al. (2021), DL has already achieved significantly better results than traditional SE techniques.
@@
However, most of the existing code learning research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models Sun et al. (2022). The challenge behind this is preparing training data for code models is heavy work. Although today developers can easily collect raw code data from the open-source community such as GitHub  [^3] https://github.com, data labeling is still expensive, especially for code-related tasks. It requires huge human effort and professional domain knowledge.
@@
To tackle this problem, data augmentation is a famous technique that increases the diversity and quality of training data based only on existing labeled data. In the code learning field techniques such as code refactoring methods and text-oriented data augmentation methods can be directly used for code augmentation. Unfortunately, existing studies Bielik and Vechev (2020); Yu et al. (2022) stated that code refactoring methods have limited advantages in the accuracy improvement of code models. Besides, the recent work Dong et al. (2025) revealed that code models trained using existing data augmentation methods still suffer from the low robustness issue. Therefore, exploring more effective and reliable code augmentation methods is an emerging task.
@@
To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision Cubuk et al. (2020); Yang et al. (2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding. Specifically, at each training epoch, GenCode first generates training candidates by using different code augmentation methods. Here, we consider different types of methods such as semantic-preserving code refactoring methods and syntax-breaking text augmentation methods. After that, GenCode selects valuable training data from the candidates with the guidance of influence scores. The influence score used in GenCode is simply defined as the loss value of code data produced by the code model. Finally, the selected samples are used for the model training. We conducted comprehensive experiments to study how GenCode can improve the accuracy and robustness of code models. Based on that, we answer the following three research questions:
RQ1:How effective is GenCode in producing accurate code models? The results demonstrate that, on average, GenCode can enhance the accuracy of pre-trained code models by up to 3.63% and 2.92% compared to the training without using data augmentation and the SOTA method MixCode, respectively. For code-specific LLMs, Gencode achieves average improvements of 0.93% for StarCoder2 and 0.84% for Qwen2.5-Coder. Notably, our statistical tests show GenCode consistently outperforms baseline methods in terms of accuracy improvement.
RQ2: How effective is GenCode in producing robust code models? The results demonstrate that GenCode produces more robust pre-trained code models compared to existing data augmentation methods. Specifically, GenCode achieves an average robustness improvement of 8.42% over models trained without data augmentation and 4.90% over models trained with MixCode. However, for code-specific LLMs, the robustness gains are modest, with improvements of less than 1%.
RQ3: How does the influence score affect the effectiveness of GenCode? The results suggest that using the influence score based on maximum loss values is the most effective strategy among the considered variants, including minimum loss value selection and random selection. Besides, even GenCode with random selection outperforms existing code augmentation methods.
RQ4: How does the data selection method affect the effectiveness of GenCode? Experimental results indicate that GenCode achieves superior performance in terms of accuracy improvement. However, the gradient-based method outperforms GenCode and other data selection methods in enhancing model robustness.
@@
  • We propose the first generation-and-selection-based framework, GenCode  [^4] https://github.com/zemingd/GenCode, which automatically enriches the training data for boosting the code understanding models.
@@
  • We empirically demonstrate that GenCode can produce more accurate and robust code understanding models compared to existing data augmentation methods.
  • We conduct an ablation study to explore how different influence scores and data selection methods affect the effectiveness of GenCode.
2 Background
In this section, we introduce the necessary knowledge that is needed to understand our work, code representation, data augmentation for code understanding, and robustness testing of code model.
2.1 Code Representation
@@
Code representation is a crucial technique that converts source code into a format interpretable by Deep Neural Networks (DNNs), facilitating effective learning of source code features. Since source code can be seen as sequences of tokens or structural graph Allamanis et al. (2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated program classification Puri et al. (2021) and code clone detection Wang et al. (2020). Existing methods Allamanis et al. (2018) for code representation can be roughly divided into two groups, sequence-based representation and graph-based representation. Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens Alon et al. (2019a). Graph-based representation learns the structural information of code. Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information Dinella et al. (2020); Yasunaga and Liang (2020).
@@
In this work, we focus on code data preprocessing and study the powerful pre-trained code models and code-specific LLMs that achieved impressive results on code understanding tasks.
2.2 Data Augmentation for Code Understanding
@@
Data augmentation is a well-studied technique in the machine learning field that increases the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort. For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym Feng et al. (2021).
@@
Recently, data augmentation has gained attention from software engineering researchers. Various data augmentation methods designed for source code have been proposed Gao et al. (2020); Yefet et al. (2020); Allamanis et al. (2021); Pour et al. (2021); Bui et al. (2021); Wang et al. (2022); Park et al. (2023); Yang et al. (2023a); Dong et al. (2024a), with the most general framework involving the use of code refactoring, e.g., local variable renaming and dead store, to create more code data without breaking the semantics of the original codes. Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information Lacerda et al. (2020).
@@
Existing works Bielik and Vechev (2020); Yu et al. (2022); de Paula Rodrigues et al. (2023); Dong et al. (2025, 2024b) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models. For instance, a recent empirical study Dong et al. (2024b) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation. Besides, the SOTA code augmentation method, MixCode, which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring. However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
@@
However, existing studies Bielik and Vechev (2020); Yu et al. (2022) have stated that these code refactoring methods have limited advantages in boosting the training of code models. To solve this problem, we propose a simple yet effective code augmentation framework that is not only unaffected by the dataset and model constraints but also applicable to both classification and non-classification tasks.
2.3 Robustness Testing of Code Model
Robustness is an important property of DL models which indicates how the DL model can handle data with different distribution of the training data. Normally, there are two types of robustness of DL models, natural robustness and adversarial robustness. Natural robustness indicates how models perform when facing the natural distribution shift, for example, data in different environments and newly collected data. Adversarial robustness refers to the ability of models to defend the adversarial attacks. Robustness testing of code models has been addressed recently. Researchers proposed distribution shifted code (e.g., codes have the same functionality but are written by different programmers) data to study the natural robustness of code models. Besides, multiple adversarial attack methods have been proposed to test the adversarial robustness of code models, for example, the recent work Yang et al. (2022) proposed to change the variable name of code to force the model to make wrong predictions.
In this work, we employ adversarial attack methods Dong et al. (2025) proposed for code to evaluate the adversarial robustness of our trained models. Following the existing work Dong et al. (2023), we construct the natural robustness datasets to evaluate the performance of all code models.
3 Definition
@@
Definition 1 (Program Code)
A program code  [^5] For simplicity, we use code in the paper. is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: contentcontent contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
Definition 2 (Code Model)
A code model is defined as a 3-tuple $M=\left(parameter,\mathcal{P},Y\right)$, where parameterparameter contains all the weights in code model $M$, $\mathcal{P}$ is the input code, and $Y$ is the label of $\mathcal{P}$ given a specific downstream code task.
Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 Wang et al. (2021).
Definition 3 (Code Augmentation)
Given a code $P$, a code augmentation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)\&\mathcal{P^{\prime}}.content\neq\mathcal{P}.content$.
Here, code augmentation methods can be any method that can change the content of program statements such as text-oriented data augmentation methods (e.g., synonym replacement and code refactoring methods (e.g., dead if statement adding).
Definition 4 (Semantic-Preserving (Breaking))
Given a code model $M$, a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, semantic-preserving (breaking) code augmentation indicates $\mathcal{P}.se=\mathcal{P^{\prime}}.se$ & $M(\mathcal{P})=M(\mathcal{P^{\prime}})$ ($\mathcal{P}.se\neq\mathcal{P^{\prime}}.se$ & $M(\mathcal{P})\neq M(\mathcal{P^{\prime}})$).
Definition 5 (Syntax-Preserving (Breaking))
@@
4 Assumption and Preliminary Study
GenCode utilizes the influence score to select the most valuable data for the code model training, where the influence score is defined as the loss value. Therefore, this approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training. To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data. Specifically, the study is conducted on the Google Code Jam (GCJ) Yang et al. (2022) dataset and CodeBERT pre-trained code model. The detailed steps are as follows:
@@ 4 Assumption and Preliminary Study
Step 4. We use the Pearson Correlation Coefficient (PCC) Schober et al. (2018) metric to check the correlation between the loss values and the accuracies.
We consider three training stages, which are the Early stage (Initialized, Epochs 0–9), the Mid stage (Middle, Epochs 10–19), and the Late stage (Well-trained, Epochs 20-29). Fig. 1 visualizes the results of our preliminary study. Here, we considered three types of code models, the initialized model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios.
(a) Initialized
(b) Middle
(c) Well-trained
Figure 1: Correlation between loss values and code model accuracy.
The results indicate that there is a positive correlation between the loss values of data and the accuracy of code models trained by using these data. From the initialized model to the well-trained model, this correlation becomes weaker but still exists. We conjecture that when the model is in the early stage, most of the data that have high loss values with more informative and lead to better model generalization. However, when training proceeds, the data that still have high loss values are the model itself cannot learn. As training progresses, the correlation between loss values and accuracy gradually weakens, which may be attributed to overfitting or the model’s limited capacity to learn from inherently difficult samples. This observation supports the hypothesis that the strength of this correlation is indicative of the model’s generalization ability. In particular, the decline in correlation from early to later training stages highlights the shifting informativeness of high-loss samples. Leveraging this insight, we design a data augmentation framework that prioritizes informative samples, especially in the early stages, to enhance the training effectiveness of code models.
@@
5 Methodology
5.1 Overview
The goal of GenCode is to produce code models with better performance on test data by training models on carefully generated data. Given a training set of program $\mathcal{P}_{train}$, its transformed version $\mathcal{P^{\prime}}_{train}$, and a set of test programs $\mathcal{P}_{test}$, after the training process $\mathsf{train}\left(\cdot\right)$ , $\varrho(M^{\prime}(\mathcal{P}_{test}))>\varrho(M(\mathcal{P}_{test}))$, where $M.parameter=\mathsf{train}(M,\mathcal{P}_{train})$ & $M^{\prime}.parameter=\mathsf{train}(M^{\prime},\mathcal{P^{\prime}}_{train})$, and $\varrho\left(\cdot\right)$ is the performance measurement.
GenCode follows the generation-and-selection workflow and utilizes different types of code augmentation methods to generate diverse code data first, then selects code data that are most useful for training as the training data. Algorithm 1 shows how GenCode works. In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14). Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data.